# Geometric Solution of the Loop Equation and Quark Confinement in Planar QCD

### v2

@@ Abstract
We present an exact analytic area law for planar QCD satisfying the MM loop equation for arbitrary loops. The solution is constructed using a quaternionic Hodge-dual minimal surface in a 16-dimensional auxiliary space. Unlike random surfaces, this construction defines a topological string theory where the action is proportional to the area of the conformal map of the 2D disk, possessing no bulk quantum fluctuations.
The Wilson loop factorizes into a confining dressing factor and the perturbative term: $W[C]=W_{pert}[C]e^{-\kappa S[C]}$. This corresponds to a non-perturbative gluon condensate vacuum supporting planar graph fluctuations. We prove the inequality $S[C]\geq\sqrt{2}A[C]$ (where $A[C]$ is the Euclidean minimal area), establishing a sufficient condition for confinement.
The area law is defined explicitly as $S[C]=2\sqrt{2}(\lambda_{3}+\lambda_{4})$ using the two largest eigenvalues of the Douglas-Gram matrix. We demonstrate that this functional is an exact additive zero mode of the loop diffusion operator, protected algebraically by the Bianchi identity. This ensures global validity for all regular loops ($C\in H^{1/2}(\mathbb{S}^{1},\mathbb{R}^{4})$).
Matching this confining factor with the gluon condensate via the QCD Operator Product Expansion (OPE) determines the physical string tension $\kappa\sim\Lambda_{QCD}^{2}$ (up to a universal numerical factor). This framework provides a first-principles derivation of confinement in large-$N_{c}$ QCD and new tools for the high-lying meson spectrum.
@@ 1 Introduction
Since the pioneering work of Ken Wilson [Wilson1974Confinement], understanding confinement in Yang–Mills theory remains one of the central challenges of modern physics. Conventional approaches based on the local action obscure the geometric structure of the Wilson loop, where confinement is expected to emerge as an area law. In this work, we use the loop–space calculus [migdal2025SQYMflow] to reformulate the problem in purely geometric terms, leading to its analytical solution. In [migdal2025SQYMflow], the general Euler-Lagrange (EL) equations for the SD/ASD surface were derived, and an exact solution was found for the class of planar loops. We construct a Hodge-dual minimal surface, such that the corresponding area law exactly solves the loop equation for the Yang–Mills gradient flow. Here, we generalize this by presenting a complete holomorphic solution for arbitrary regular loops (with Fourier coefficients decreasing faster than $1/n$). This holomorphic solution supersedes the planar case constructed in Ref. [migdal2025SQYMflow], extending it to arbitrary loops and simultaneously correcting a typographical error $Y^{A}_{\mu}=0$ in the proof of the planar solution in that paper.
@@ 2 Basic definitions and duality equations
With $(\ast\Sigma)_{\mu\nu}=\tfrac{1}{2}\epsilon_{\mu\nu\rho\sigma}\Sigma_{\rho\sigma}$, the SD/ASD parts are $\Sigma^{(\pm)}_{\mu\nu}=\tfrac{1}{2}(\Sigma_{\mu\nu}\pm\ast\Sigma_{\mu\nu})$, satisfying $\ast\Sigma^{(\pm)}_{\mu\nu}=\pm\Sigma^{(\pm)}_{\mu\nu}$. The Euler-Lagrange variational equations for this functional were written down in [migdal2025SQYMflow]. These nonlinear differential equations were solved in that paper only for a special case of planar boundary loop $C$. In this paper, we generalize that solution for an arbitrary loop.
@@
$\frac{\delta W}{\delta\sigma_{\mu\nu}(\theta)}=$
$\frac{\delta}{\delta C^{\prime}_{\mu}(\theta-0)}\frac{\delta W}{\delta C^{\prime}_{\nu}(\theta+0)}-\frac{\delta}{\delta C^{\prime}_{\mu}(\theta+0)}\frac{\delta W}{\delta C^{\prime}_{\nu}(\theta-0)}$ (2.3)
@@
The basic observation made in that paper was there is a SD solution similar to the Yang–Mills multi-instanton. The loop space diffusion equation (2.5) is identically satisfied for any functional $S[C]$ with SD or ASD area derivative
@@
The last equation (Bianchi identity) is valid at the kinematical level as a Jacobi identity for a triple commutator (see [migdal2025SQYMflow] for a proof). The matrix in the Dirichlet boundary conditions (BC) for $X\sim QC$ should be chosen so that the area derivative represents SD or ASD tensor. This would make the exponential of the Hodge-dual minimal area an exact solution to the loop equation.
@@ 3 Geometric Definition of the Physical Area
Finally, this maximized minimal projection area depends on the parametrization of the bounding curve, unlike the Wilson loop which is reparametrization invariant by definition. This last ambiguity is eliminated in the same way as in the Euclidean minimal surface: by minimizing the resulting functional over the group of orientation-preserving diffeomorphisms of the circle, $C(\theta)\to C(\phi(\theta))$ with $\phi^{\prime}>0$. Combining these steps, the full definition for a loop decomposable into sub-loops $C_{i}$ is:
@@ 4 A Self-Dual Holomorphic Ansatz
We propose a minimal ansatz that satisfies the self-duality condition (4.2) using two independent holomorphic functions $f_{1}(z)$ and $f_{2}(z)$. Let the only non-zero components of $U$ be:
@@
The EL equations of this functional are satisfied by Cauchy-Riemann equations for $f_{1},f_{2}$, so that this functional reduces to the boundary data. By the Dirichlet identity, $S_{\chi}=2\sqrt{2}\sum_{n\geq 1}n\big(|\hat{f}_{1,n}|^{2}+|\hat{f}_{2,n}|^{2}\big)$. We now fix the boundary conditions by relating $f_{1},f_{2}$ to the boundary loop $C_{\mu}$ in a rotated basis:
@@
Throughout we assume the boundary loop $C:S^{1}\to\mathbb{R}^{4}$ is regular in the sense that the spectral expansion converges in the Douglas–Gram matrix, so that the Dirichlet identity leading to (4.6) is valid, and that the Fourier/Hilbert-transform manipulations used in this paper are justified. This corresponds to $C\in H^{\tfrac{1}{2}}(S^{1},\mathbb{R}^{4})$
@@ 5 Invariant Area and the Spectral Law
The true, physical $O(4)$-invariant area $S_{\mathrm{phys}}$ was defined as the maximum area that can be generated by our functional form over all possible orientations $R$ of the loop $C$. For this holomorphic solution
@@
Let $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$ be the eigenvalues of $G$. The maximum in (5.1) is evaluated by first applying the cyclic property of the trace, which gives $\max_{R\in O(4)}2\sqrt{2}\,\operatorname{tr}((R^{T}PR)G)$. We can define a new projector $P^{\prime}=R^{T}PR$. Since $P$ is a fixed rank-2 projector, $P^{\prime}$ is also a rank-2 projector. The group $O(4)$ acts transitively on the space of all 2-planes, $\operatorname{Gr}(2,4)$. This means that by maximizing over all $R\in O(4)$, the projector $P^{\prime}$ sweeps through the entire set of all possible rank-2 projectors. Therefore, our maximization is mathematically identical to $2\sqrt{2}\max_{P^{\prime}\in\operatorname{Gr}(2,4)}\operatorname{tr}(P^{\prime}G)$. By the Ky Fan theorem [Horn:2012:MatrixAnalysis:2nd], this maximum is precisely the sum of the two largest eigenvalues, $\lambda_{3}+\lambda_{4}$.
@@ 5.1 Projector Alignment
The physical area $S_{phys}$ is realized by maximizing the base functional $S_{\chi}$ over the global $\mathrm{O}(4)$ rotation group $\mathbf{R}$: $S_{phys}=\max_{\mathbf{R}\in\mathrm{O}(4)}2\sqrt{2}\operatorname{tr}(\mathbf{P}(\mathbf{RGR}^{\top}))$. The initial fixed projector $\mathbf{P}$ is defined by the orthonormal basis vectors $\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ associated with the Hodge parity $\chi=\pm 1$:
$\mathbf{u}_{1}$ $=\frac{1}{\sqrt{2}}(\mathbf{e}_{1}+\mathbf{e}_{2})$ (5.3)
$\mathbf{u}_{2}$ $=\frac{1}{\sqrt{2}}(\mathbf{e}_{3}+\chi\mathbf{e}_{4})$ (5.4)
The optimal rotation $\mathbf{R}_{max}=\mathbf{R}_{2}\mathbf{R}_{1}$ maps this initial SD basis onto the optimal basis $\{\mathbf{v}_{3},\mathbf{v}_{4}\}$ corresponding to the largest eigenvalues of $\mathbf{G}$, $\lambda_{3}$ and $\lambda_{4}$.
5.1.1 1. Projection and Orthogonalization (the combined operation $\mathbf{R}_{1}$
The first operation $\mathbf{R}_{1}$ starts with alignement of the initial plane $\operatorname{span}(\mathbf{u}_{1},\mathbf{u}_{2})$ with the optimal eigenspace $\mathcal{V}_{34}=\operatorname{span}(\mathbf{v}_{3},\mathbf{v}_{4})$. This is achieved by projecting $\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ orthogonal to the null space spanned by the two lowest eigenvectors $\{\mathbf{v}_{1},\mathbf{v}_{2}\}$. Let $\mathbf{P}_{12}=\mathbf{v}_{1}\mathbf{v}_{1}^{\top}+\mathbf{v}_{2}\mathbf{v}_{2}^{\top}$ be the projector onto this null space. The intermediate vectors $\mathbf{u}^{\prime}_{i}$ are given by: $\mathbf{u}^{\prime}_{i}=\operatorname{Normalize}\left[\mathbf{u}_{i}-\mathbf{P}_{12}\cdot\mathbf{u}_{i}\right]$
The second part of this operation $\mathbf{R}_{1}$ orthogonalizes these two vectors: $\mathbf{u}^{\prime\prime}_{2}=\operatorname{Normalize}\left[\mathbf{u}_{2}-(\mathbf{u}_{2}\cdot\mathbf{u}_{1})\mathbf{u}_{1}\right]$ After these two transformations, the basis vectors $\mathbf{u}_{1},\mathbf{u}_{2}$ become orthogonal basis in the Eigenspace ($\mathbf{R}_{2}$. Therefore, this combined transformation is an $O(4)$ rotation.
5.1.2 2. Alignment within the Eigenspace ($\mathbf{R}_{2}$)
The second rotation $\mathbf{R}_{2}$ performs an $\mathrm{O}(2)$ transformation within the optimal plane $\mathcal{V}_{34}$ to fix the alignment of the resulting basis. This step aligns the first intermediate vector $\mathbf{u}^{\prime}_{1}$ with the highest eigenvector $\mathbf{v}_{3}$: $\mathbf{R}_{2}:\quad\text{An }\mathrm{O}(2)\text{ rotation in }\mathcal{V}_{34}\text{ such that }\mathbf{R}_{2}\mathbf{u}^{\prime}_{1}\propto\mathbf{v}_{3}$ The combined rotation $\mathbf{R}_{max}$ defines the required boundary condition $\mathbf{X}=\mathbf{Q}\cdot\mathbf{R}_{max}\mathbf{C}$ for the minimal surface, which yields the spectral area $S_{phys}$.

### v3

# Geometric Solution of the Loop Equation and Quark Confinement in QCD

@@ Abstract
We present an exact analytic confining area law for pure $\mathrm{SU}(N_{c})$ Yang–Mills theory, satisfying the hierarchy of multiloop equations for arbitrary regular loops and an arbitrary number of colors $N_{c}$. The solution is constructed using a quaternionic Hodge-dual minimal surface in a 16-dimensional auxiliary space. Unlike random surfaces, this construction defines an effectively topological string theory in which the bulk dynamics reduces to the area of a conformal map of the two-dimensional disk, with all nontrivial dynamics encoded in the boundary loop.
The expectation value of a Wilson loop factorizes into a confining dressing factor and a perturbative term,
$W[C]=W_{\text{pert}}[C]\,e^{-\kappa S[C]}.$
We show that the geometric factor $e^{-\kappa S[C]}$ is an exact multiplicative zero mode of the loop-space diffusion operator which satisfies the full MM multiloop hierarchy for any finite $N_{c}$. The same confining area law applies equally to the planar limit and to the physical case $N_{c}=3$. The functional $S[C]$ obeys the inequality $S[C]\geq\sqrt{2}\,A[C]$ (where $A[C]$ is the Euclidean minimal area), which provides a sufficient condition for confinement.
The area law is given explicitly by the spectral formula:
$S[C]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$
in terms of the two largest eigenvalues of the Douglas–Gram matrix of the loop. We demonstrate that this functional is an exact additive zero mode of the loop diffusion operator, protected algebraically by the loop-space Bianchi identity, and that it is globally defined for all regular loops $C\in H^{1/2}(\mathbb{S}^{1},\mathbb{R}^{4})$. Matching this confining factor with the gluon condensate via the QCD operator product expansion (OPE) determines the physical string tension, $\kappa\sim\Lambda_{\text{QCD}}^{2}$ up to a universal numerical factor. This framework yields an analytic, first-principles derivation of an area law for Wilson loops in QCD and provides new tools for computing the high-lying meson spectrum.
@@ 1 Introduction
Since the pioneering work of Ken Wilson [Wilson1974Confinement], understanding confinement in Yang–Mills theory remains one of the central challenges of modern physics. In the Wilson loop formulation, confinement is expected to emerge as an area law for large loops. However, in the usual local Lagrangian description, the geometric structure of the loop is largely hidden. An appealing alternative is to reformulate the dynamics directly in loop space, where the MM equations Ref. [MM1981NPB, Mig83] provide a nonlocal, but exact, description of pure $\mathrm{SU}(N_{c})$ Yang–Mills theory in terms of Wilson loop functionals.
In this work, we adopt the loop–space calculus developed in Ref. [migdal2025SQYMflow] to recast the confinement problem in purely geometric terms and to solve the resulting loop equation analytically. In Ref. [migdal2025SQYMflow] the general Euler–Lagrange (EL) equations for a self–dual/anti–self–dual (SD/ASD) Hodge–dual area functional were derived in an auxiliary quaternionic target space, and an exact solution was found for the restricted class of planar loops. The corresponding Hodge–dual minimal surface yields an area law that exactly solves the loop equation for the Yang–Mills gradient flow, but only for loops whose Douglas–Gram matrix has rank two. Moreover, the proof of the planar solution in [migdal2025SQYMflow] contained an error ($Y^{A}_{\mu}=0$), which we correct here. We validate the claim of that paper for planar loops, and generalize it for arbitrary loops.
The first main result of the present paper is a complete holomorphic construction of the quaternionic Hodge–dual minimal surface for arbitrary regular loops $C\in H^{1/2}(\mathbb{S}^{1},\mathbb{R}^{4})$ (equivalently, loops with Fourier coefficients decreasing faster than $1/n$). We show that the minimized Hodge–dual area can be written in a simple spectral form:
$S[C]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4}),$
where $\lambda_{i}$ are the ordered eigenvalues of the Douglas–Gram matrix Ref. [Douglas1931] associated with the loop. This geometric functional is defined by a min–max–min principle (over bulk embeddings, internal isometries, and boundary reparametrizations), is additive under canonical splitting at self–intersections, and has an SD/ASD area derivative protected by the loop–space Bianchi identity. As a consequence, $S[C]$ is an exact zero mode of the loop–space diffusion operator, and the multiplicative dressing
$W[C]\;\longrightarrow\;W[C]\;e^{-\kappa S[C]},$
preserves the loop equation for the Yang–Mills gradient flow for any value of the parameter $\kappa$.
The second main result is that this geometric dressing actually extends beyond the planar approximation: the same functional $S[C]$ defines an exact multiplicative zero mode of the full MM multiloop hierarchy for an arbitrary number of colors $N_{c}$. The key ingredients here are the strict additivity of $S[C]$ at all intersection points and the fact that the same loop operator acts on every member of the hierarchy. When we apply the dressing
$W_{n}(C_{1},\dots,C_{n})\;=\;$
$W_{n,\mathrm{pert}}(C_{1},\dots,C_{n})\,\exp\left(-\kappa\sum_{i=1}^{n}S[C_{i}]\right)$ (1.1)
to the chain of multiloop equations, the geometric factor factors out of each term, cancels between the two sides of the equations, and leaves the entire hierarchy invariant for any finite $N_{c}$. Thus, the confining area law originally obtained in the planar limit is in fact a universal feature of the pure Yang-Mills vacuum, equally valid for the physical case $N_{c}=3$.
A further consequence is a rigorous confinement theorem. Comparing our spectral functional with the Euclidean minimal area $A[C]$ (or, equivalently, with the trace of the Douglas–Gram matrix), we prove the inequality
$S[C]\;\geq\;\sqrt{2}\,A[C],$
which provides a universal lower bound on the effective string tension associated with any regular loop. Finally, by matching the small–loop expansion of the dressed Wilson loop to the operator product expansion (OPE) in QCD, we relate the parameter $\kappa$ to the gluon condensate and the QCD scale, $\kappa\sim\Lambda_{\mathrm{QCD}}^{2}$ up to a universal numerical factor. In this way, the geometric Hodge–dual minimal surface furnishes an exact confining area–law solution of the loop equations in pure $\mathrm{SU}(N_{c})$ Yang–Mills theory and a quantitative bridge to the computation of the hadron spectrum.
@@ 2 Basic definitions and duality equations
With $(\ast\Sigma)_{\mu\nu}=\tfrac{1}{2}\epsilon_{\mu\nu\rho\sigma}\Sigma_{\rho\sigma}$, the SD/ASD parts are $\Sigma^{(\pm)}_{\mu\nu}=\tfrac{1}{2}(\Sigma_{\mu\nu}\pm\ast\Sigma_{\mu\nu})$, satisfying $\ast\Sigma^{(\pm)}_{\mu\nu}=\pm\Sigma^{(\pm)}_{\mu\nu}$. The Euler-Lagrange variational equations for this functional were written down in [migdal2025SQYMflow]. These nonlinear differential equations were solved in that paper only for a special case of a planar boundary loop $C$. In this paper, we generalize that solution for an arbitrary loop.
@@
$\frac{\delta W}{\delta\sigma_{\mu\nu}(\theta)}=\frac{\delta^{2}W}{\delta C^{\prime}_{\mu}(\theta-0)\delta C^{\prime}_{\nu}(\theta+0)}-\mu\leftrightarrow\nu$ (2.3)
@@
The basic observation made in that paper was that there is an SD solution similar to the Yang–Mills multi-instanton. The loop space diffusion equation (2.5) is identically satisfied for any functional $S[C]$ with SD or ASD area derivative
@@
The last equation (Bianchi identity) is valid at the kinematical level as a Jacobi identity for a triple commutator (see [migdal2025SQYMflow] for a proof). The matrix in the Dirichlet boundary conditions (BC) for $X\sim QC$ should be chosen so that the area derivative represents the SD or ASD tensor. This choice would make the exponential of the Hodge-dual minimal area an exact solution to the loop equation.
@@ 3 Geometric Definition of the Physical Area
Finally, the maximized minimal projection area depends on the parametrization of the bounding curve, unlike the Wilson loop, which is reparametrization-invariant by definition. This last ambiguity is eliminated in the same way as in the Euclidean minimal surface: by minimizing the resulting functional over the group of orientation-preserving diffeomorphisms of the circle, $C(\theta)\to C(\phi(\theta))$ with $\phi^{\prime}>0$. Combining these steps, the full definition for a loop decomposable into sub-loops $C_{i}$ is:
@@ 4 A Self-Dual Holomorphic Ansatz
We propose a minimal ansatz that satisfies the self-duality condition (4.2) using two independent holomorphic functions $f_{1}(z)$ and $f_{2}(z)$. Let the only nonzero components of $U$ be:
@@
The EL equations of this functional are satisfied by the Cauchy-Riemann equations for $f_{1},f_{2}$, so that this functional reduces to the boundary data. By the Dirichlet identity, $S_{\chi}=2\sqrt{2}\sum_{n\geq 1}n\big(|\hat{f}_{1,n}|^{2}+|\hat{f}_{2,n}|^{2}\big)$. We now fix the boundary conditions by relating $f_{1},f_{2}$ to the boundary loop $C_{\mu}$ in a rotated basis:
@@
Throughout, we assume the boundary loop $C:S^{1}\to\mathbb{R}^{4}$ is regular in the sense that the spectral expansion converges in the Douglas–Gram matrix, so that the Dirichlet identity leading to (4.6) is valid, and that the Fourier/Hilbert-transform manipulations used in this paper are justified. This corresponds to $C\in H^{\tfrac{1}{2}}(S^{1},\mathbb{R}^{4})$
@@ 5 Invariant Area and the Spectral Law
The true, physical $O(4)$-invariant area $S_{\mathrm{phys}}$ was defined as the maximum area that can be generated by our functional with all possible rotations $R$ of the loop $C$. For this holomorphic solution
@@
Let $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$ be the eigenvalues of $G$. The maximum in (5.1) is evaluated by first applying the cyclic property of the trace, which gives $\max_{R\in O(4)}2\sqrt{2}\,\operatorname{tr}((R^{T}PR)G)$. We can define a new projector $P^{\prime}=R^{T}PR$. Since $P$ is a fixed rank-2 projector, $P^{\prime}$ is also a rank-2 projector. The group $O(4)$ acts transitively on the space of all 2-planes, $\operatorname{Gr}(2,4)$. This transitivity means that by maximizing over all $R\in O(4)$, the projector $P^{\prime}$ sweeps through the entire set of all possible rank-2 projectors. Therefore, our maximization is mathematically identical to $2\sqrt{2}\max_{P^{\prime}\in\operatorname{Gr}(2,4)}\operatorname{tr}(P^{\prime}G)$. By the Ky Fan theorem [Horn:2012:MatrixAnalysis:2nd], this maximum is precisely the sum of the two largest eigenvalues, $\lambda_{3}+\lambda_{4}$.
@@ 5.1 Projector Alignment
The physical area $S_{phys}$ is realized by maximizing the base functional $S_{\chi}$ over the global $\mathrm{O}(4)$ rotation group. The maximum is achieved when the boundary projector aligns with the principal axes of the loop’s geometry:
$S_{phys}=\max_{\mathbf{R}\in\mathrm{O}(4)}2\sqrt{2}\operatorname{tr}\left(\mathbf{P}\mathbf{R}\mathbf{G}\mathbf{R}^{\top}\right).$ (5.3)
The initial fixed projector $\mathbf{P}$ is defined by the orthonormal basis $U=(\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3},\mathbf{u}_{4})$ associated with the Hodge parity $\chi=\pm 1$. Explicitly, the basis vectors are:
$\mathbf{u}_{1}$ $=\tfrac{1}{\sqrt{2}}(\mathbf{e}_{1}-\mathbf{e}_{2}),$ $\mathbf{u}_{2}$ $=\tfrac{1}{\sqrt{2}}(\mathbf{e}_{3}-\chi\mathbf{e}_{4}),$
$\mathbf{u}_{3}$ $=\tfrac{1}{\sqrt{2}}(\mathbf{e}_{1}+\mathbf{e}_{2}),$ $\mathbf{u}_{4}$ $=\tfrac{1}{\sqrt{2}}(\mathbf{e}_{3}+\chi\mathbf{e}_{4}).$ (5.4)
In this basis, the projector takes the diagonal form:
$\mathbf{P}=U\operatorname{diag}(0,0,1,1)U^{\top}.$ (5.5)
Similarly, the Douglas-Gram matrix $\mathbf{G}$ is diagonalized by the orthogonal matrix $V$ formed by its eigenvectors $\{\mathbf{v}_{i}\}$ sorted by the magnitude of their eigenvalues $\lambda_{i}$:
$\mathbf{G}=V\operatorname{diag}(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})V^{\top}.$ (5.6)
The optimal rotation $\mathbf{R}_{\ast}$ is the orthogonal transformation that maps the eigenbasis of $\mathbf{G}$ directly onto the basis of $\mathbf{P}$. This rotation matrix is given by the ratio of the two basis matrices:
$\mathbf{R}_{\ast}=UV^{\top}.$ (5.7)
Substituting this rotation into the trace functional yields the maximal value directly as the dot product of the ordered spectra:
$\operatorname{tr}(\mathbf{P}\mathbf{R}_{\ast}\mathbf{G}\mathbf{R}_{\ast}^{\top})$
$=\operatorname{tr}\left((U\Lambda_{P}U^{\top})(UV^{\top})(V\Lambda_{G}V^{\top})(VU^{\top})\right)$
$=\operatorname{tr}\left(\Lambda_{P}\Lambda_{G}\right)=\lambda_{3}+\lambda_{4}.$ (5.8)